OpenStudy (anonymous):

i need precal help! -1/2+log5/3(4x+3)=1/2 -1+log5/3(2x+1/3)=-3

OpenStudy (campbell_st):

so can I check, the 1st question is \[-\frac{1}{2} + \log (\frac{5}{3} (4x + 3) )= \frac{1}{2}\]

OpenStudy (campbell_st):

what is the base of the log..?

OpenStudy (anonymous):

the base is 5/3, it seems odd because ive never seen a log w a fraction as b

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... well then its quite easy add 1/2 to both sides and it becomes \[\log_{\frac{5}{3} }(4x + 3) = 1\] now you need log laws for powers and bases \[if ~~~~\log_{a}(x) = b ~~~then~~~~ a^{\log_{a}(x)} = a^b ~~or~~~ x = a^b\] so by raising both sides of the equation to powers of 5/3 you get \[\frac{5}{3}^{\log_{\frac{5}{3}}(4x +3)} = (\frac{5}{3})^1\] which becomes \[4x + 3 = \frac{5}{3}\] now solve for x hope it helps

OpenStudy (anonymous):

@campbell_st thank you so much!